Question:1

It is claimed that an automobile is driven in the average less than 20,000 km /year. To test this claim, a random sample of 100 auto-mobiles owners are asked to keep a record of the kilometers, they travel. Would you agree with this claim if the random sample showed an average of 23,500 kilometers and a standard deviation of 3900 kilometers. Use a 0.01 level of significance.


Question: 2

Test the hypothesis that the average content of containers of a particular lubricant is 10 liters if the contents of a random sample of 10 containers are 10.2, 9.7, 10.1, 10.3, 10.1, 9.8, 9.9, 10.4, 10.3, and 9.8 litters.

Use a 0.01 level of significance and assume that the distribution is Normal.


1
Expert's answer
2022-05-30T03:54:12-0400

1. The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge20000"

"H_1:\mu<20000"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=99" and the critical value for a left-tailed test is "t_c =-2.364606."

The rejection region for this left-tailed test is "R = \{t:t<-2.364606\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{23500-20000}{3900/\sqrt{100}}=8.9744"

Since it is observed that "t=8.9744>-2.364606=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for left-tailed, "df=99" degrees of freedom, "t=8.9744" is "p=1," and since "p=1>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is less than 20000, at the "\alpha = 0.01" significance level.


2. Mean  

"\dfrac{1}{10}(10.2+9.7+10.1+10.3+10.1"




"+9.8+9.9+10.4+10.3+9)=10.06"

Variance



"s^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n-1}=\dfrac{1}{10-1}(0.0196+0.1296"




"+0.0016+0.0576+0.0016+0.0676+0.0256"

"+0.1156+0.0576+0.0676)=\dfrac{0.544}{9}"


"s=\sqrt{s^2}=\sqrt{\dfrac{0.544}{9}}\approx0.246"


The following null and alternative hypotheses need to be tested:

"H_0:\mu=10"

"H_1:\mu\not=10"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=9" and the critical value for a two-tailed test is "t_c =3.249823."

The rejection region for this two-tailed test is "R = \{t:|t|>3.249823\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{10.06-10}{0.246/\sqrt{10}}=0.7713"

Since it is observed that "|t|=0.7713<3.249823=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed, "df=9" degrees of freedom, "t=0.7713" is "p=0.460299," and since "p=0.460299>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 10, at the "\alpha = 0.01" significance level.